Which is greater: 
 (1/2) of (6/7) or (2/3) of (3/7)......

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (1/2) × (6/7) and (2/3) × (3/7)
By the rule Multiplication of fraction,
Product of fraction = (product of numerator)/ (product of denominator)
Then
= (1/2) × (6/7)
= (1 × 6)/ (2 × 7)
= (1 × 3)/ (1 × 7)
= (3/7) ......(i)
And,
= (2/3) × (3/7)
= (2 × 3)/ (3 × 7)
= (2 × 1)/ (1 × 7)
= (2/7) …...(ii)
By comparing [i] and [ii],
Clearly,
(3/7) < (2/7)

Thanks